Digital Test
Hewlett-Packard 524C

Manufacturer: Hewlett-Packard
Model: 524C
Year: 1961
Schematic: RE,MF
Max. Rate: 10 MHz
Display: NIXIE
Digits: 8
Logic: tube

The 'C' version of the HP 524. The 524C uses NIXIE tubes for the display instead of the 1-of-10 neon-bulb displays of the other 524 versions. The unit I have was manufactured in 1961, however the schematics in the 524C/D manual are dated 1959, making this a quite early instrument to use NIXIEs.

The 1962 HP catalog presents a list price of $2400.

The photo shows a unit with a 525A 100MC Frequency Converter plugged into the front panel.

This unit had an exasperating fault which caused the count to reset to zero immediately at the end of the gate period. One would watch it speed through the count and then have no opportunity to observe the result. After tracking the problem to a subtle issue around the gate circuitry, the unit is fully functional.

For more about the 524 models see the HP 520 Series.

The 524C removed from it's cabinet. The window cover over the NIXIEs has also been removed. 526B Time Interval plug-in and 525A 100MC Frequency Converter plug-in also present.

The circuitry of the 524 is organised onto three decks: the rear, right side and bottom. The tube side of all three decks faces the inside, while the component/wiring side is always around the outside.

Looking inside through the front panel plug-in hole at the tube side of the decks. It is possible (with some difficulty) to access all the tubes through this hole without taking the chassis out of the cabinet.

Top view showing AC-4 counter string.

Rear deck: Power Supply on the right and Timebase on the left.

Right-side deck: Units Counter.

This gives an idea of the complexity required to get the first decade to count at 10Mhz.

Bottom deck: Gate and Control.

Note the two massive power supply transformers at the bottom.

Looking through the left side with the tubes lit up. Makes a nice room heater. The bright light is from the decimal point lamp.

- Unit Log -

Serial Number: 116-03509
Year of Manufacture: 1961 (AC-4K counter module stamped with "6 28 61", capacitor with "4-61", coil with "5-61", etc.)
Other Identification: Stock No. 6625-473-5411
Model K33-999 Serial 224
Contract No. AF 33 (604) 31304
U.S. Property
Date of Receipt: 1994 Dec
State upon Receipt: Functional. Reset/cycling does not work properly until after a long warmup (15-30 min). Internals very clean, a little grubby on the front panel. Cabinet, which should be the light pinky-brown HP used in the 1950s, has been spray-painted silver.

V1101 in trigger unit has a 12AU7 installed, should be a 5963.

AC-4K in MSD position is a later version than all the others, dated Feb 66, so would appear to be a replacement.

With 525A plug-in (100MHz Frequency Converter, Serial No. 048-10743).

Date: 1994 Dec
Procedure: Reverse-engineered. Decaying power cord replaced.

Date: ~2002
Procedure: Photocopied HP manual received. Pages sorted and placed in binder.

Date: ~2003
Procedure: 526B plug-in (Time Interval, Serial No. E024-00177) received. Has metal plate with "Department of Transport, Telecommunications and Electronics Branch, Serial No. RFC 64006".

Date: 2006 May 16
Symptoms: LSD displays "6" after reset. Gate lamp does not light, even with toggling of manual gate switch.

Date: 2008 Mar 09
Symptoms: As on 2006 May 16.
Procedure: Power supply voltages checked. Everything is low, pointing to the C-195 which measures around 135V. 160K R in voltage divider for C- measures around 200 ohms. Doing some brushing around the leads of the adjustment pot solves this. Must have been some contamination causing a short to the pot housing.

Unit now back to usual state of taking a 1/2 hour before it will hold the count after the gate period. Also a problem with timebase divider chain, operation stops at the 1 second gate period.

Date: 2008 Mar 09
Procedure: 526B Time Interval plug-in finally tested. Functional.

Date: 2008 Mar 10
Symptoms: Timebase gate operation stops at 1 second setting.
Procedure: Plate V246.5 is only 15V. Cathode V245.1 shorted by a piece of hair-thin wire caught on the terminal board. Where it came from is a mystery, blown in by the fan perhaps. It was probably a similar piece producing the C- problem of Mar 09.

Date: 2008 Mar 11
Symptoms: Old problem of not holding the count (resetting to 0) at end of gate period. 100KC & 10MC Check work on 0.001 gate period, 100KC check works on 0.01 gate period.
Analysis: In period mode with external variable oscillator, unit functions properly for periods shorter than ~1.7 mS; longer periods result in reset behaviour.

Output of reset blocking oscillator (grey lead of transformer) observed to propagate out a second pulse coincident with the stop edge of the gate when the gate pulse-width goes beyond the pulse-width of the blocking oscillator.

There is some overshoot on the stop edge which may be due to a slow clamp diode which appears to be the second trigger. Components around blocking oscillator seem to be good: bias voltage for blocking oscillator is -17.4V.

Capacitance of scope lead connected to green lead of transformer coil observed to fix the problem. Not clear if this is just covering an overshoot problem from the Start-Stop FF (although the overshoot appears to be normal in that it appears on both outputs) or whether the transformer may have changed in characteristics. Second pulse seems quite unstable viewed on scope.

Several 5963's tried in V208, and swapping V206 and 207 (6AH6), to no benefit.

Solution: Addition of a 100pF capacitor to ground with clip leads works, but fails after soldered in with shorter leads. 180pF narrows the failure range to approx. 1.8mS - 2.2mS, where second pulse rides on the trailing edge of the main pulse. 390pf seems to fix it.

Note: soldering around 82pF cap and diode in grid circuit of blocking oscillator appears to have been previously worked post-manufacture.

Date: 2008 Mar 14
Procedure: V1101, which had a 12AU7 installed, replaced with a proper 5963.

Date: 2008 Mar 14
Procedure: Tubes and tube decks cleaned, some AC-4 counter string tubes remain to be done. V257 5Y3 has broken guide in bakelite base: glued back on.

Date: 2008 Mar 15
Procedure: 4th from right DP lamp was $/hash$44, replaced with $/hash$47 (all others are $/hash$47).

Date: 2008 Mar 17
Symptoms: 10MC check not consistent across range of gate periods, varies by count of 0-20 in lowest digits.
Analysis: Guess is that gate time is now sloppy due to additional capacitance in plate circuit V208.6 (see Mar 11).
Solution: 390pF capacitor from Mar 11 removed, 27pF capacitor to ground installed directly on grid V208.2. Now stable ±1 across all ranges on both 100KC and 10MC checks. Unit now fully functional.

Date: 2008 Mar 19
Procedure: Another attempt to isolate the reset problem. 27 pf cap from Mar 17 removed. Observed that a very small capacitance to ground on the output of the blocking oscillator (grey lead) will also solve the problem. Suspicion falls on C219 (39pF), but replacement does not solve problem. Increasing C219 to 51pF improves the situation, but too much (~90pF) results in complete loss of the reset pulse. Various attempts to bypass existing caps around the reset switch and the blocking oscillator with a good cap made to see if one is failing and thus leading to a loss of C in the circuit. No help.

27pF cap reinstalled as per Mar 17, as it seems to be the most reliable fix.

  .. | Hewlett-Packard 524C | ..
Early Digital Test Equipment